# dragapart.github.io

### v1

@@ 3.3 Generalization to Real Data via Domain Randomization
Most objects in GAPartNet [18], which we use to construct the dataset used to train our model (see Sec. 4), have synthetic textures not representative of real-world objects. Since these textures usually lack variability in color, models trained on these renderings can potentially ‘cheat’ by generating the same pixel values everywhere without an implicit understanding of where the pixel comes from in the original image $y$. Empirically, we find training the model jointly on a version of the renderings in which each object part is assigned a random color, as shown in Fig. 1 and Fig. 3b and sometimes called domain randomization [66], significantly improves the model’s performance on real-world images. The effectiveness of this training strategy is validated in Sec. 5.
@@ 3.4 Downstream Applications
While our primary goal is to demonstrate that an image generator can learn a motion prior for articulated objects at part-level, we also explore the potential of our DragAPart in several downstream applications. By operating at the level of parts, our approach is complementary to prior works that focus on the task of repositioning [75, 68] and can address different applications than them.
@@
First, DragAPart can be used to segment the moving parts of an object. Several prior works [2, 74, 65, 29] have used pre-trained generators for 2D segmentation, but focus on part semantic rather than part mobility. We suggest that DragAPart, as a motion prior, can be used to segment the part that would move in response to a drag (useful, e.g., to predict affordances).
In order to segment mobile parts, we follow [2] and extract internal features from the forward part of the denoiser $\Phi$. However, this work, which looks at semantics, extracts a single set of features from the input image $x$. Instead, because our goal is to obtain the segmentation of a part that moves prompted by a drag, we run the forward pass twice, one with the drag conditioning and one without, and compute the difference between the resulting features. This difference represents the effect of the drag in the diffusion model’s internal representation.
@@
In order to estimate the parts’ types and parameters using DragAPart, we start by rendering $K$ images $y_{k}$ of the object $\mathcal{M}$ from random cameras $C_{k}$. We also project the 3D drag to pixel space $d_{\text{pixel}}^{k}\in\Omega\times\mathbb{Z}^{2}$. Then, we use DragAPart to generate images $x_{k}\sim\mathbb{P}(x_{k}|y_{k},\left\{d_{\text{pixel}}^{k}\right\})$ of the dragged object under $C_{k}$. Following [36], using the generated images, we estimate the motion parameters $(t,p_{\text{motion}})$ by minimizing the objective function
@@
where $R$ is the rendering function and $T$ transforms the moving parts based on the motion estimation. While [36] solves Eq. 2 using backpropagation, we find gradient-based optimization prone to local optima, leading to suboptimal estimation. Since the search space is relatively low dimensional, we optimize Eq. 2 directly using grid search.
@@ 4 The Drag-a-Move Dataset
Key to training our model is a suitable dataset with drag annotations. The dataset must thus contain a collection of triplets $(x,y,\mathcal{D})$ where image $y$ represents the initial state of the object, $\mathcal{D}$ the collection of drags applied to $y$, and image $x$ the same object in a new state compatible with the action of the drags. Moreover, we wish the state change to be minimal, in the sense of affecting only the smallest subset of object parts that still result in a physically plausible state.
Creating such a dataset using real-world images would require carefully capturing photos of an articulated object in two articulation states, while ensuring consistency in camera position, lighting conditions and background across captures, and then manually annotating or reconstructing the drag. Fortunately, recent works on new-view synthesis indicated that generator models can be fine-tuned effectively using synthetic data only. We thus opt to create a synthetic dataset, which allows us to control the articulation and lighting conditions, and to obtain the ground-truth drags directly from the 3D model.
We build our dataset on an existing 3D synthetic dataset, GAPartNet [18], which adds rich, part-level annotations to the objects of [73, 38]. By rendering the assets under different articulation, we can obtain synthetic data with desired properties, as described next.
@@
Among all the objects in GAPartNet, we select those that have at least one part which is annotated with one of the labels “Hinge Handle”, “Slider Drawer”, “Hinge Lid” and “Hinge Door”, which results in $763$ candidate 3D models spanning $16$ categories. See Sec. 0.A.3 for details.
@@
The annotation of each object contains its kinematic tree structure, where nodes are rigid parts and edges are joints (fixed or not). Our goal is to sample a diversity of starting and ending states (before and after drag) of the different parts, including cases in which only a subset of parts move. We generate with equal probability two kinds of animations. In the first, some parts are fixed and locked to their “closed” or “rest” state, whereas the other parts transition continuously between two extreme states (e.g., a drawer fully closed or fully open, Fig. 3 right). The second kind is similar, but the stationary parts are set to a random state, and the moving parts transition between two randomly selected states (Fig. 3 left). This is further detailed in Sec. 0.A.3.
@@
Since we have ground-truth 3D data with articulation annotation, we can project the motion $(p,q)\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ of any visible 3D point on the moving part back to pixel space to obtain a drag $(u,v)$. However, this would result in a dense set of drags, akin to optical flow, whereas we are interested in only selecting one drag per part. In order to generate a sparse subset of drags, for each moving part we first sample a part in the subtree rooted at that part uniformly at random. Then, we sample a point on the visible surface of a subpart with probabilities proportional to its displacement due to the parent part motion. We use the trajectory of that point as the drag for the moving part.
@@
For each 3D asset, we render $48$ articulation animations, with 36 frames each, under random lighting conditions and camera viewpoints by means of the SAPIEN [73] renderer. The images and drags are constructed at resolution $512\times 512$. We also render a version of the dataset where each part of the 3D model is given a random, monochromatic color texture (Fig. 3 right), which we find useful for the model’s generalizability (Sec. 3.3).
@@ 5 Experiments
We quantitatively and qualitatively show that our method outperforms prior works, and provide extensive ablation studies to validate our design choices. Please refer to Appendix 0.A and Appendix 0.B for additional results and details.
@@ 5.1 Datasets and Implementation Details
We assess the DragAPart on the test split of our Drag-a-Move dataset (Sec. 4), which contains $23$ object assets not seen during training across $10$ categories and $150$ image pairs. Furthermore, we also conduct evaluations on real images sourced from the Common Object in 3D (CO3D) dataset [52] and the Amazon Berkeley Objects (ABO) dataset [11]. For these, we use SAM [31] for object segmentation and manually define plausible drags.
@@
We train a separate model for humans on the Human3.6M dataset [25]. Following previous work [5, 10], we use actors S1, S5, S6, S7 and S8 for training and S9 and S11 for testing. We randomly sample two frames, as the conditioning image and target, within a video sequence. The corresponding drags are derived using the annotated paths of keypoints that show sufficient displacement.
@@
We explore two LDM architectures, SD [53] and DiT [46]. In our implementation, we minimally alter the original architecture to enable a more accurate comparison of the capacities of both architectures. Detailed modifications and our training schedule are provided in Sec. 0.A.1 and Sec. 0.A.2.
@@ 5.2 Comparisons with Prior Work
We perform a thorough comparison of DragAPart to the state-of-the-art generative models which are applicable to this task: iPoke [5], DragDiffusion [59], DragonDiffusion [42], DragNUWA [75], and Motion Guidance [17]. For DragDiffusion, DragonDiffusion and Motion Guidance, we use the official implementations and checkpoints because they are training-free methods. Since Motion Guidance is designed for dense motion manipulation, we include two versions in our comparison, one given a dense motion field as input and the other adapted to take only sparse drags. We train iPoke using the publicly available code on our Drag-a-Move dataset and use its released checkpoint trained on Human3.6M [25]. For DragNUWA, we use the released checkpoint. As it is a large-scale open-domain video generator, we do not fine-tune it with our data. iPoke and DragNUWA generate videos, from which we take the last frame for comparison. Additionally, we compare with another naive baseline, InstructPix2Pix [6], which fine-tunes SD [53] on a dataset of paired images and corresponding text edits generated with a large language model [7]. As it is conditioned on textual instructions, we manually craft prompts to convey the manipulation of the drags.
Refer to captiona)b)c)d)e)0123456InputInstructPix2PixDragDiffusionDragonDiffusionDragNUWAMotion GuidanceOursInputiPokeDragDiffusionDragonDiffusionDragNUWAOursGround TruthInputInstructPix2PixiPokeMG(S)MG(D)OursGround Truth
@@
Quantitative Comparisons.
We conduct quantitative comparisons on the test split of I.D. Drag-a-Move and Human3.6M in Tab. 1. DragAPart significantly outperforms the state-of-the-art methods on all standard metrics, including pixel-level PSNR, patch-level SSIM [67], and feature-level LPIPS [77]. This includes iPoke [5], which has been trained with the identical experimental setting as DragAPart, and DragNUWA, which is a large pre-trained generative model with motion control, but that mainly focuses on moving objects as a whole, but does not understand the motion resulting from part-level interactions. InstructPix2Pix and Motion Guidance are excluded in quantitative comparisons because curating a textual instruction for every image pair in the test set is impractical, and Motion Guidance takes an hour to generate a single image, making it infeasible to run it for every test example in our dataset.
Qualitative Comparisons.
The qualitative results are shown in Fig. 4. As can be seen, despite our best efforts, InstructPix2Pix is never able to deform the object significantly, which motivates using dragging for conditioning instead of language. Methods that build on Stable Diffusion [53] latent features (DragDiffusion, DragonDiffusion, Motion Guidance) tend to directly copy the appearance from the handle points to the target points without moving the part as a whole (a-d3, e4). For DragDiffusion and DragNUWA, the object is sometimes re-scaled (a2, b2, c4) or repositioned (a4, c2), rather than performing a part-level movement. For Motion Guidance, even when the dense motion field is given (e4), it fails to preserve the object identity very well (a-b5, e4). By contrast, by using a data-driven approach, our model learns to generate images that preserve faithfully the identity and that exhibit physically-plausible motion. Applied to humans (d), our method produces a result that is closest to the ground truth, without changing the identity of the actor, as opposed to iPoke (d1).
@@
Table 2: Ablations on Architecture and Flow Encoding. The flow encoding we proposed in Sec. 3.2, when injected in every LDM block at different resolutions, enhances the model’s performance, regardless of the underlying LDM architecture.
@@ 5.3 Ablation Analysis and Limitations
We present ablation studies on the two key components of DragAPart: the multi-resolution flow encoding (Sec. 3.2) and training with random texture (Sec. 3.3).
As shown in Tab. 2, it is not sufficient to inject the drag conditioning through the input layer only (“Input Only” in Tab. 2). As previous works [75, 68] have discovered, injecting this motion conditioning at every LDM block (“Every Block” in Tab. 2) is essential. Our multi-resolution drag encoding better enforces disentanglement among different part motion than using a convolution network (Fig. 5 left). It outperforms the previous method and the improvement generalizes to both LDM architectures (Tab. 2).
The experimental results also validate the importance of our domain randomization strategy (Sec. 3.3). Without training on the random texture renderings, the model’s performance on out-of-distribution data drops significantly (Tab. 3), and generalizes poorly to unseen categories (Fig. 5 right).
@@
5.4 Qualitative Results on Real Data
Fig. 6 presents more real images generated by our model. Our model is capable of preserving fine-grained texture details (a), generating reasonable shades (b), handling thin structures (c), compositing multi-part motion (d), “dreaming” up internal structures of the object (e), and generalizing to categories not seen during training (f). More results are presented in Sec. 0.B.1.
Refer to captionInputGenerationInputGenerationInputGenerationa)b)c)d)e)f)
@@ 5.5 Application Results
Fig. 7 (left) shows two motion analysis results, where we accurately recover the motion type and parameters specified by a drag. This also indicates that our DragAPart model is relatively consistent across different viewpoints. In Fig. 7 (right) we show that DragAPart’s implicit part-level understanding can be leveraged to produce a coarse segmentation of moving parts. Notably, the segmentation faithfully reflects the input drag. In Sec. 0.B.3, we provide a quantitative comparison of the segmentation quality on I.D. Drag-a-Move (with available ground-truth masks) to baselines against CLIP [50], DINO [9, 44] and SD [53].
@@ 6 Conclusion
We presented DragAPart, an image generator that uses drags as an interface for part-level motion interaction. By using a small amount of synthetic data, which we also contribute, and domain randomization DragAPart can respond to drags by interpreting them as fine-grained part-level interactions with the underlying object. Partly due to a new drag encoder, we have obtained better results than other methods on this task, including generalizing to real data. We have also demonstrated some applications of the motion model that DragAPart provides.
Acknowledgments. We would like to thank Minghao Chen, Junyu Xie and Laurynas Karazija for insightful discussions. This work is in part supported by a Toshiba Research Studentship and ERC-CoG UNION 101001212.

### v2

@@ 3.3 Generalization to Real Data via Domain Randomization
Most objects in GAPartNet [17], which we use to construct the dataset used to train our model (see Sec. 4), have synthetic textures not representative of real-world objects. Since these textures usually lack variability in color, models trained on these renderings can potentially “cheat” by generating the same pixel values everywhere without an implicit understanding of where the pixel comes from in the original image $y$. Empirically, we find training the model jointly on a version of the renderings in which each object part is assigned a random color as shown in Fig. 1 and Fig. 3 (right), sometimes called domain randomization [tobin17domain], significantly improves the model’s performance on real-world images. The effectiveness of this training strategy is validated in Sec. 5.
@@ 3.4 Downstream Applications
While our primary goal is to demonstrate that an image generator can learn a motion prior for articulated objects at part-level, we also explore the potential of our DragAPart in several downstream applications. By operating at the level of object parts, our approach is complementary to prior works that focus on the task of repositioning [72, 65] and can address different applications than them.
@@
First, DragAPart can be used to segment the moving parts of an object. Several prior works [2, 71, 63, 27] have used pre-trained generators for 2D segmentation, but focus on part semantic rather than part mobility. We suggest that DragAPart, as a motion prior, can be used to segment the part that would move in response to a drag (useful, e.g., for predicting affordances).
In order to segment mobile parts, we follow [2] and extract internal features from the forward pass of the denoiser $\Phi$. However, [2] looks at semantics and only extracts a single set of features from the input image. Instead, because our goal is to obtain the segmentation of the part that moves prompted by a drag, we run the forward pass twice, once with the drag conditioning and once without, and compute the difference between the resulting features. This difference represents the effect of the drag in the diffusion model’s internal representation.
@@
In order to estimate the parts’ motion types and parameters using DragAPart, we start by rendering $K$ images $y_{k}$ of the object $\mathcal{M}$ from random cameras $C_{k}$. We also project the 3D drag to pixel space $d_{\text{pixel}}^{k}\in\Omega\times\mathbb{Z}^{2}$. Then, we use DragAPart to generate images $x_{k}\sim\mathbb{P}(x_{k}|y_{k},\left\{d_{\text{pixel}}^{k}\right\})$ of the dragged object under $C_{k}$. Following [34], using the generated images, we estimate the motion parameters $(t,p_{\text{motion}})$ by minimizing the objective function
@@
where $R$ is the rendering function and $T$ transforms the moving parts based on the motion estimation. While [34] solves Eq. 2 using back-propagation, we find gradient-based optimization prone to local optima, leading to sub-optimal estimation. Since the search space has relatively low dimensionality4447777 for revolute type and $4$ for prismatic type., we optimize Eq. 2 directly using grid search.
@@ 4 The Drag-a-Move Dataset
Key to training our model is a suitable dataset with drag annotations. It must thus contain a collection of triplets $(x,y,\mathcal{D})$, where image $y$ represents the initial state of the object, $\mathcal{D}$ is the collection of drags applied to $y$ and image $x$ denotes the same object in a new state compatible with the action of the drags. Moreover, we wish the state change to be minimal, in the sense that each drag only affects the smallest subset of object parts that still result in a physically plausible state.
Creating such a dataset using real-world images would require carefully capturing photos of an articulated object in two articulation states while ensuring consistency in camera position, lighting condition and background across captures, and then manually annotating or reconstructing the drags. Fortunately, recent works on novel-view synthesis have indicated that generative models can be effectively fine-tuned using synthetic data only. We thus choose to create a synthetic dataset, which allows us to control the articulation and lighting conditions, and to obtain the ground-truth drags directly from the 3D model.
We build our dataset on an existing 3D synthetic dataset, GAPartNet [17], which adds rich, part-level annotations to the objects of [70, 36]. By rendering the assets under different articulations, we can obtain synthetic data with desired properties, as described next.
@@
Among all the objects in GAPartNet, we select those that have at least one part which is annotated with one of the labels “Hinge Handle”, “Slider Drawer”, “Hinge Lid” and “Hinge Door”, which results in $763$ candidate 3D models spanning $16$ categories. See the sup. mat. for details.
@@
The annotation of each object contains its kinematic tree structure, where nodes are rigid parts and edges are joints (fixed or not). Our goal is to sample a diversity of starting and ending states (before and after drag) of the different parts, including cases in which only a subset of parts move. We generate with equal probability two kinds of animations. In the first, some joints are fixed and locked to their “closed” or “rest” state, whereas the other joints transition continuously between two extreme states (e.g., a drawer fully closed or fully open, Fig. 3 right). The second kind is similar, but the stationary joints are set to a random state, and the moving joints transition between two randomly selected states (Fig. 3 left). This is further detailed in the sup. mat.
@@
Since we have ground-truth 3D data with articulation annotation, we can project the motion $(p,q)\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ of any visible 3D point on the moving part back to pixel space to obtain a drag $(u,v)$. However, this would result in a dense set of drags, akin to optical flow, whereas we are interested in only selecting one drag per part. In order to generate a sparse subset of drags, for each moving joint we first sample a part in the subtree (of the entire kinematic structure) rooted at that joint uniformly at random. Then, we sample a point on the visible surface of that part with probabilities proportional to its displacement due to the parent joint’s motion. We use the trajectory of that point as the drag for the moving parts. This process is fast, enabling us to generate different drag combinations on the fly during training to augment the dataset further.
@@
For each 3D asset, we render $48$ articulation animations, with 36 frames each, under random lighting conditions and camera viewpoints by means of the SAPIEN [70] renderer. This amounts to $40$ million unique image pairs $(x,y)$ for training purposes. We also render a version of the dataset where each part of the 3D model is given a random, monochromatic color texture (Fig. 3 right), which we find useful for the model’s generalizability (Sec. 3.3).
@@ 5 Experiments
We show that our method outperforms prior works both quantitatively and qualitatively, and conduct ablation studies to validate our design choices.
@@ 5.1 Datasets and Implementation Details
We assess DragAPart on the test split of our Drag-a-Move dataset (Sec. 4), which contains $150$ image pairs from $23$ objects not seen during training spanning $10$ categories. Furthermore, we also conduct evaluations on real images sourced from the Common Object in 3D (CO3D) dataset [50] and the Amazon Berkeley Objects (ABO) dataset [11]. For these, we use SAM [29] for object segmentation and manually define plausible drags.
@@
We train a separate model for humans on the Human3.6M dataset [23]. Following previous works [5, 10], we use actors S1, S5, S6, S7 and S8 for training and S9 and S11 for testing. We randomly sample two frames, as the conditioning image and target, within a video sequence. The corresponding drags are derived using the annotated paths of keypoints that show sufficient displacement.
@@
We evaluate SD [51] and DiT [44] as LDM architectures. In our implementation, we minimally alter the original architecture to enable a more accurate comparison of the capacities of both architectures. Detailed modifications and our training schedule are provided in the sup. mat.
@@ 5.2 Comparisons with Prior Work
We perform a thorough comparison of DragAPart to the state-of-the-art generative models which are applicable to this task: iPoke [5], DragDiffusion [57], DragonDiffusion [40], DragNUWA [72], and Motion Guidance [16]. For the training-free methods (DragDiffusion, DragonDiffusion and Motion Guidance), we use their official implementations. Since Motion Guidance is designed for dense motion manipulation, we include two versions in our comparison, one given a dense motion field as input and the other adapted to take only sparse drags. We train iPoke using the publicly available code on our Drag-a-Move dataset and use its released checkpoint trained on Human3.6M [23]. For DragNUWA, we use the released checkpoint. As it is a large-scale open-domain video generator, we do not fine-tune it with our data. iPoke and DragNUWA generate videos, from which we take the last frame for comparison. Additionally, we compare with another naive baseline, InstructPix2Pix [6], which fine-tunes SD [51] on a dataset of paired images and corresponding text edits generated with a large language model [7]. As it is conditioned on textual instructions, we manually craft prompts to convey the manipulation of the drags.
a)b)c)d)e)0123456InputInstructPix2PixDragDiffusionDragonDiffusionDragNUWAMotion GuidanceOursInputiPokeDragDiffusionDragonDiffusionDragNUWAOursGround TruthInputInstructPix2PixiPokeMG(S)MG(D)OursGround Truth
@@
Quantitative Comparisons
are conducted on the test split of I.D. Drag-a-Move and Human3.6M in Tab. 1. DragAPart significantly outperforms the state-of-the-art methods on all standard metrics, including pixel-level PSNR, patch-level SSIM, and feature-level LPIPS. This includes iPoke [5], which has been trained with the identical experimental setting as DragAPart, and DragNUWA, which is a large pre-trained generative model with motion control, but mainly focuses on moving objects as a whole and does not yield motion resulting from part-level interactions. InstructPix2Pix and Motion Guidance are excluded in quantitative comparisons because curating a textual instruction for every image pair in the test set is impractical, and we could not run Motion Guidance on the entire test set as it requires one hour per instance.
Qualitative Comparisons
are shown in Fig. 4. Despite our best efforts, InstructPix2Pix is never able to deform the object significantly, which motivates using dragging for conditioning instead of language. Methods that build on Stable Diffusion [51] latent features (DragDiffusion, DragonDiffusion, Motion Guidance) tend to copy the appearance of the handle points to the target points without moving the entire part (a-d3, e4). For DragDiffusion and DragNUWA, the object is re-scaled (a2, b2, c4) or repositioned (a4, c2), rather than performing a part-level motion. For Motion Guidance, even when the dense motion field is given (e4), it fails to preserve the object identity very well (a-b5, e4). By contrast, by using a data-driven approach, our model learns to generate images that preserve faithfully the identity and that exhibit physically plausible motion. Applied to humans (d), our method produces a result that is closest to the ground truth, without changing the identity of the actor, as opposed to iPoke (d1).
Qualitative Results on Real Data.
Fig. 6 shows real images generated by our model, which can (a) preserve fine-grained texture details, (b) generate reasonable shades, (c) handle thin structures, (d) compose multi-part motion, (e) “dream up” the internal structures of the object, and (f) generalize to categories not seen during training. More results are presented in the sup. mat.
@@
Table 2: Ablations on Architecture and Flow Encoding. The flow encoding we proposed in Sec. 3.2, when injected into every LDM block at different resolutions, enhances the model’s performance, regardless of the underlying LDM architecture.
@@ 5.3 Ablation Analysis and Limitations
We present ablation studies on the two key components of DragAPart: the multi-resolution flow encoding (Sec. 3.2) and training with random texture (Sec. 3.3). As shown in Tab. 2, it is not sufficient to inject the drag conditioning through the input layer only (“Input Only” in Tab. 2). As previous works [72, 65] have discovered, injecting this motion conditioning at every LDM block (“Every Block” in Tab. 2) is essential. Our multi-resolution drag encoding better enforces disentanglement among different part motion than using a convolution network (Fig. 5 left). It outperforms the previous methods and the improvement generalizes to both LDM architectures (Tab. 2). Note also the importance of our domain randomization strategy (Sec. 3.3). Without training on the random texture renderings, the model’s performance on out-of-distribution data drops significantly (Tab. 3), and generalizes poorly to unseen categories (Fig. 5 right).
@@
InputGenerationInputGenerationInputGenerationa)b)c)d)e)f)
@@ 5.4 Application Results
Fig. 7 (left) shows two motion analysis results, where we accurately recover the motion type and parameters specified by a drag. This also indicates that our DragAPart model is relatively consistent across different viewpoints. In Fig. 7 (right) we show that DragAPart’s implicit part-level understanding can be leveraged to produce a coarse segmentation of moving parts. Notably, the segmentation faithfully reflects the input drag. In the sup. mat., we provide a quantitative comparison of the segmentation quality on I.D. Drag-a-Move (with available ground-truth masks) against baselines using CLIP [48], DINO [9, 42] and SD [51].
@@ 6 Conclusion
We presented DragAPart, an image generator that uses drags as an interface for part-level dynamics. By using a small amount of synthetic data, which we contributed, and domain randomization, DragAPart can respond to drags by interpreting them as fine-grained part-level interactions with the underlying object. Partly thanks to a new drag encoder, we have obtained better results than other methods on this task. The model generalizes well to real-world images and novel categories. We have also demonstrated some applications of the motion prior that DragAPart provides.
Acknowledgments.
We thank Minghao Chen, Junyu Xie and Laurynas Karazija for insightful discussions and our sponsors Toshiba, ERC-CoG UNION 101001212 and EPSRC SYN3D EP/Z001811/1.